Friday, May 30, 2008

Incoming Internet E-Mail

We have been experiencing some issues within incoming mail and are presently working with tech support to resolve the matter.  E-mail is being delivered but is backlogged several hours.

Thursday, May 22, 2008

Lansing-Reilly & CHP network bounce

Between 8:00am and 8:20am this morning, a network repair was done for parts of Lansing-Reilly and the CHP building. Computers in these two buildings were unable to reach the network, including the internet and Email.

All repairs have been completed, and all services should be back to normal.

Tuesday, May 20, 2008

Internet E-mail

Incoming e-mail is backed up once again with approximately 60,000 messages that should be delivered in the next couple of hours.

Monday, May 19, 2008

Internet E-mail Delivery

We are presently backed up to about 4:30 pm.  Mail is processing and should be caught up within the next couple of hours.


Sunday, May 18, 2008

Incoming Internet E-Mail

There have been a number of external problems in regard to receiving incoming e-mail.  These issues are persisting but mail is processing through.  There is a slight backlog that should be processed through in the next two hours.

Friday, May 16, 2008

Blackboard update 05/16 8:00 p.m. - Online

Blackboard services are available to users. Thank you for your patience during this outage.

Blackboard update 05/16 5:00 p.m

Blackboard will be available shortly.

Thursday, May 15, 2008

BlackBoard Update - 5:15pm

The drive rebuild is complete, however, the original server is unusable. The data for both the application and the database has been recovered from the failed system and is currently being migrated to the second server. This migration of data is expected to take several more hours.

Streaming Server Maintenance Completed

Maintenance on has been completed. All streaming services have been restored.

Blackboard Update at 10:00 am

The network team worked through the night to rebuild the bad drive.  The process requires a significant amount of time and is still running.  While this is occurring, a plan B has been put into action to bring up the second server and restore all data to it.  We anticipate it will be several more hours before this process is complete.


Wednesday, May 14, 2008

Blackboard Still Down

Work continues to restore services.


Update on Blackboard status 05/14 5:10 p.m.

Blackboard outage is caused by hardware failure. Procedure to preserve integrity of content is running at this time. Time needed for this procedure is approximately 2 hours.

Knowledge/Blackboard Server Down

The server is temporarily down due to a hard drive issue.  Work will continue until it is back up.


Weblink Available 9:30 AM on 5/14/08

The physical move of the Weblink server is complete. The Weblink application is available as of 9:30 AM. Thank you, ITS Dept.

Dental server issue

The Dental server UDM-TN-DENTAL was rebooted at 8:50am because of memory pool problems. Our users could not get to any network resources on that server. The reboot will clear up the issue, and we will investigate into what caused the memory leak.

Streaming Server Maintenance will be off line from 10:00 a.m. to 10:30 a.m Thursday, May 15th for maintenance. All streaming services will be unavailable during this time.

Tuesday, May 13, 2008

UHC + Corktown Connectivity

ATT has repaired the T1 lines that were down. We are back to full capacity.

Weblink Down May 14, 2008 8:30 - 9:30 AM

Weblink will be unavailable tomorrow morning (May 14, 2008) from 8:30 until 9:30 AM for the purpose of physically moving the server to a new rack.
Thank you,
ITS Department

Academic Lab Server Temporarily Down

The server is temporarily down for relocation and will be up in less than 30 minutes.

UHC + Corktown Connectivity

There is currently an effort by ATT to restore 2(two) T1 Lines. One of these lines provides Data to UHC - the other provides partial service to the phone system at Corktown. Unless the phone traffic at Corktown increases substantially users may not even notice a problem. However, the Data at UHC is offline at present. Both problem are being handled by ATT and should be repaired soon. We will update when the work has been completed. Thank You.

Sunday, May 04, 2008

Dorm Internet Access

Late this evening it was discovered there were problems with the connection for the resident students. This has been resolved in the past 15 minutes. If there are further issues with the dorm access, please respond to this post.

Saturday, May 03, 2008

Incoming Internet E-Mail

Incoming internet e-mail is backed up as of about 2:00 am.  The mail is processing through and should be caught up by 10:00 am.